OpenStudy (anonymous):

Help! Will fan and medal! Assume that 1a_1+2a_2+...+na_n=1, where the a_j are real numbers. As a function of n, what is the minimum value of 1a_1^2+2a_2^2+...+na_n^2?

OpenStudy (kindaljoshconall):

@Bill_Cipher

jhonyy9 (jhonyy9):

so this mean that 1a_1 =1 2a_2 =2*2 =4 3a_3=3*3 =9 4a_4 =4*4 = 16 ..... na_n =na_n = n^2 for example for n=2 1a_1 =1*1 =1 2a_2 = 2*2 =2 1a_1 +2a_2 = 1 => 1+2=3=1 1a_1^2 = 1^2 = 1 2a_2^2 = 2*2^2 =8 1a_1^2 +2a_2^2 =9 =3^2 = 1^2 = 1 so from this result that 1a_1^2 +2a_2^2 +... +na_n^2 = 1

OpenStudy (anonymous):

1 isn't correct though. (I tried already.)

jhonyy9 (jhonyy9):

@ganeshie8 please your opinion about this - thank you in advance

OpenStudy (deer_slayer13):

2

OpenStudy (anonymous):

How did you get that? BTW it isn't right.

OpenStudy (vishweshshrimali5):

Basically we have... \[\sum_{i = 1}^{n} i*a_i = 1\] And you want to find out minimum value of \[\sum_{i = 1}^{n}i*a_i ^2\]

OpenStudy (vishweshshrimali5):

Hmm... well I may have something in mind..

OpenStudy (vishweshshrimali5):

Have you read about Cauchy Schwarz inequality?

OpenStudy (anonymous):

Yes.

OpenStudy (vishweshshrimali5):

Or let me write it like this instead of that... \[\large{|\sum\sqrt{i}(\sqrt{i}*{a_i})|^2 \le \sum|\sqrt{i}^2| \sum|(\sqrt{i}a_i)^2|}\]

OpenStudy (vishweshshrimali5):

Yeps!! This is much better :D

OpenStudy (vishweshshrimali5):

Now, I know the LHS value... we can find out \(\sum_{i = 1}^{n} i\) and we want to find out the value of the other summation

OpenStudy (vishweshshrimali5):

Can you do after this?

OpenStudy (anonymous):

I'm not sure.

OpenStudy (vishweshshrimali5):

See... we have this formula: \(\sum_{i=1}^{n} i = \frac{n(n+1)}{2}\)

OpenStudy (anonymous):

Okay.

OpenStudy (vishweshshrimali5):

Also, LHS of the inequality will become 1^2 = 1... So, we can separate out the last summation as: \[\large{\sum_{i=1}^{n} i a_i^2 \ge \cfrac{2}{n(n+1)}}\]

OpenStudy (vishweshshrimali5):

From here you can find out the min. value